Hanna:

Line L passes through the points (a, 2a) and (b, 2b), where a ≠ b. Determine the slope of line L. A) 1 B) 2 C) -2 D) a - b E) undefined

MARC:

formula for slope is \(slope,m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\)

MARC:

let say \(point(a,2a)=(x_{1},y_{1})\) and \(point(b,2b)=(x_{2},y_{2})\)

MARC:

substitute those values into the slope formula.

MARC:

\(slope,m=\frac{2b-2a}{b-a}\)

MARC:

Factorise the 2

MARC:

\(=\frac{\color{red}{2}(b-a)}{b-a}\)

MARC:

Then,simplify it.

MARC:

\(=\frac{2}{1}\times\frac{b-a}{b-a}\)

MARC:

\(=2\times1\)

MARC:

=?

MARC:

solve it @Hanna